# A Reduction-Driven Local Search for the Generalized Independent Set Problem

### v1

@@ 6 Computational Experiments
This section is dedicated to a comprehensive evaluation of the proposed reduction rules and heuristic approaches. In order to evaluate the algorithm performance, we compare both solution quality and runtime between RLS and other benchmark algorithms; to evaluate the effect of reduction rules, we compare the algorithms with reductions and without reductions; we also conduct an ablation study to evaluate the effect of each reduction rule.
@@ 6.1 Datasets
The proposed reduction rules and the RLS algorithm are evaluated on three sets and a total of $278=216+47+15$ instances, covering various application scenarios.
SET-1. The first set of instances consists of $216$ standard instances with up to $400$ vertices and up to $71,820$ edges. These instances were first introduced by Colombi et al. (2017) and tested in Nogueira et al. (2021), Hosseinian and Butenko (2019), Zheng et al. (2024). This set of instances consists of 12 graphs with 125 to 400 vertices and 6963 to 71820 edges (available at https://or-dii.unibs.it/instances/Instances_GISP.zip).
@@
SET-3. To further evaluate the algorithms on large instances and diverse real-world scenarios, we introduce 15 instances with up to 18 million vertices and 261 million edges. The first four smaller instances represent ecosystem and road/facility networks, while the remaining 11 larger instances are drawn from hyperlink, citation, and social networks. The instances were generated using the same method for generating SET-2 (available at https://github.com/PlutoAiyi/RLS/tree/main).
@@ 6.2 Experimental Settings
All experiments are conducted on a machine with an Intel(R) Xeon(R) Platinum 8360Y CPU @ 2.40GHz and 2T RAM.
Benchmark Algorithms. To evaluate the effectiveness of our proposed reduction rules and RLS algorithm, we compare them with the best-performing algorithms for the GIS problem.
@@
  • ALS: The adaptive multi-neighborhood local search algorithm proposed by Zheng et al. (2024)
@@
The source codes of ILS-VND, LA-B&B and ALS were made available by the authors. All source codes are available at https://github.com/PlutoAiyi/RLS.
Parameter Settings. The RLS algorithm requires two parameters: the search depth $L$ and the perturbation parameter $\epsilon\in(0,1)$. We set $L=10|V|$ where $|V|$ is the number of vertices of the input graph and $\epsilon=0.2$. This setting is obtained using the general IRACE automatic parameter configuration tool (López-Ibáñez et al. 2016). For large instances like soc-pokec, LiveJ and uk2002, we decrease the search depth as $L=0.5|V|$ to encourage explorations within the cutoff time.
@@ 6.3 Performance Comparison with Existing Algorithms
We first compared five algorithms on the SET-1 instances. The exact algorithms CB&B and LA-B&B ran with a cutoff time of 3 hours (10800 seconds) for each instance. If an exact algorithm did not finish within the cutoff time, then we reported the best solution found. The remaining three heuristic algorithms ran 10 trials with a cutoff time of 30 seconds for each instance. We reported the largest value among the best-known objective values obtained in these trials. The setting is consistent with that in (Hosseinian and Butenko 2019, Nogueira et al. 2021).
In Table 2, we summarize the computational results on SET-1. In the last two columns, we show the number of instances where the corresponding algorithm performs best (#Best) and worst (#Worst) among all the five algorithms, in terms of the best known objective value and runtime, respectively. Note that the runtime of RLS refers to the wall-clock time, including both pre-processing and local search. We observe that, in all instances, the best-known objective values obtained by local search algorithms are comparable to those obtained by exact algorithms. In particular, if an exact algorithm achieves the optimal solution within 3 hours, local search algorithms can also find an optimal solution (although without proof). However, when neither of the exact algorithms finds the optimal solution, local search algorithms always produce a better one. Besides, the exact LA-B&B dominates CB&B whereas the heuristic ALS and RLS dominate ILS-VND in terms of both solution quality and runtime. Therefore, we only consider LA-B&B, ALS, and RLS for even larger instances in SET-2 and SET-3. Detailed results on this set are left in Table LABEL:tab:_Set1_detail in Appendix B.
@@
bio-yeast 1,458 980 968 66,545 300 68,572 9 68,574 $<$1 0 $<$1
soc-wiki-Vote 889 727 2,187 35,973 300 37,358 8 37,358 $<$1 0 $<$1
vc-exact_131 2,980 1,191 4,169 107,812 300 109,392 14 109,641 $<$1 0 $<$1
web-edu 3,031 4,813 1,661 109,856 300 113,781 16 114,140 $<$1 98 $<$1
MANN_a81 3,321 3,118 3,362 114,616 300 116,907 14 117,327 $<$1 0 $<$1
tech-routers-rf 2,113 1,539 5,093 93,690 300 97,799 11 97,799 $<$1 0 $<$1
vc-exact_039 6,795 2,574 8,046 254,941 300 258,137 9 259,542 $<$1 0 $<$1
ca-GrQc 4,158 3,179 10,243 167,846 300 174,654 15 174,782 11 282 $<$1
vc-exact_038 786 10,321 3,703 10,371 300 12,014 12 12,015 4 697 $<$1
vc-exact_078 11,349 4289 13,450 102,312 300 435,412 $<$1 438,671 $<$1 0 $<$1
vc-exact_087 13,590 5126 16,114 71,804 300 517,500 2 518,500 $<$1 0 $<$1
vc-exact_107 6,402 15,646 5,594 88,602 300 400,223 1 407,878 $<$1 0 $<$1
vc-exact_151 15,783 12,144 12,519 54,114 300 522,462 1 530,138 $<$1 0 $<$1
vc-exact_167 15,783 18,235 6,428 55,190 300 457,272 2 465,178 $<$1 0 $<$1
bio-dmela 7,393 18,897 6,672 280,897 300 302,449 10 302,992 $<$1 44 $<$1
vc-exact_011 9,877 19,102 6,871 190,729 300 301,789 7 303,793 $<$1 172 $<$1
web-spam 4,767 8,089 28,386 180,446 300 190,344 14 190,459 13 487 $<$1
vc-exact_026 6,140 27,159 9,608 177,898 300 201,316 17 201,503 $<$1 300 $<$1
vc-exact_001 6,160 19,655 20,552 190,729 300 215,668 13 215,941 $<$1 134 $<$1
vc-exact_024 7,620 11,367 35,926 226,540 300 235,440 14 236,018 12 1,255 $<$1
C1000.9 1,000 11,989 37,432 6,102 300 7,229 12 7,237 10 1,000 $<$1
tech-WHOIS 7,476 42,181 14,762 309,903 300 325,327 14 325,581 13 338 $<$1
p-hat700-3 700 30,398 31,242 2,937 300 4,473 $<$1 4,473 $<$1 700 $<$1
vc-exact_008 7,537 35,678 37,155 71,804 300 246,771 14 247,091 $<$1 574 $<$1
keller5 776 17,800 56,910 2,788 300 3,877 1 3,877 2 776 $<$1
vc-exact_194 1,150 39,520 41,331 4,273 300 5,419 $<$1 5,419 $<$1 1,150 $<$1
hamming10-4 1,024 66,369 23,231 2,871 300 3,660 $<$1 3,660 $<$1 1,024 $<$1
brock800_2 800 26,629 84,805 1,996 300 2,473 $<$1 2,473 $<$1 800 $<$1
brock800_4 800 26,831 85,126 1,968 300 2,501 $<$1 2,501 $<$1 800 $<$1
ca-HepPh 11,204 57,721 59,898 153,907 300 370,779 6 372,079 17 2,165 $<$1
p-hat700-2 700 90,788 32,134 2127 300 2,971 $<$1 2,971 $<$1 700 $<$1
vc-exact_196 1534 93,273 32,809 4273 300 5,321 $<$1 5,321 $<$1 1,534 $<$1
p-hat700-1 700 135,804 47,847 844 124 844 $<$1 844 $<$1 700 $<$1
ca-AstroPh 17,903 47,135 149,837 44,296 300 602,320 2 603,945 14 4,985 1
C2000.9 2,000 147,548 51,920 5477 300 6396 3 6396 3 2,000 $<$1
DSJC1000.5 1,000 59,694 189,980 1492 300 1861 1 1861 $<$1 1,000 $<$1
socfb-MIT 6,402 123,064 128,166 110,677 300 129,807 17 129,872 16 5,494 $<$1
p-hat1500-3 1,500 66,206 210,800 3856 300 7,501 14 7,501 16 1,500 $<$1
socfb-UCSB37 14,917 356,918 125,297 219,965 300 254,965 16 255,898 22 13,655 3
socfb-Duke14 9,885 374,731 131,706 125,940 300 151,169 15 151,326 18 9,003 4
p-hat1500-2 1,500 411,372 143,918 2,444 300 4,253 $<$1 4,253 $<$1 1,500 $<$1
socfb-Stanford3 11,586 136,823 431,486 215,381 300 251,350 12 251,833 14 9,380 2
socfb-UConn 17,206 448,038 156,829 170,548 300 274,637 12 275,034 20 16,025 5
p-hat1500-1 1,500 411,744 427,583 1,121 300 1,287 $<$1 1,287 1 1,500 $<$1
C2000.5 2,000 489,894 509,270 1,570 300 1,849 2 1,849 6 2,000 1
keller6 3,361 759,816 266,766 3,545 300 5,236 10 5,236 16 3,361 $<$1
C4000.5 4,000 1,957,884 2,039,848 1,601 300 1,948 20 1,950 22 4,000 9
# of best / # of instances 1/47 18/47 47/47
average running time(s) 296.3 7.9 5.9 1.4
@@
foodweb-wet 128 1,509 566 2,464 54.84 2,464 $<$1 2,464 $<$1 125 $<$1
foodweb-dry 128 1,571 535 2,226 27.87 2,226 $<$1 2,226 $<$1 125 $<$1
USAir97 332 1,589 537 9,507 2000 10,308 $<$1 10,308 $<$1 0 $<$1
powergrid 4,941 1,727 4,867 190,537 2000 193,095 59 193,339 $<$1 0 $<$1
CondMat 23,133 46,769 46,670 95,130 2000 705,901 14 711,334 45 1,648 $<$1
Email 265,009 91,093 273,388 OOM OOM 13,054,427 1 0 1
Epinion 75,879 202,653 203,087 7,400 2000 3,038,533 18 3,052,059 1 253 1
Dblp 317,080 787,268 262,598 OOM OOM 9,734,653 64 11,851 11
cnr-2000 325,557 685,686 2,053,283 OOM OOM 13,368,670 482 59,052 399
WikiTalk 2,394,385 3,494,873 1,164,692 OOM OOM 118,708,326 333 8 333
BerkStan 685,230 1,662,833 4,986,637 OOM OOM 26,031,704 1112 114,191 1031
As-Skitter 1,696,415 5,550,661 5,544,637 OOM OOM 66,582,254 723 31,173 636
soc-pokec 1,632,803 11,152,945 11,149,019 OOM OOM 43,999,402 934 881,319 134
LiveJ 4,846,609 10,710,579 32,140,658 OOM OOM 168,431,056 904 401,307 469
uk2002 18,483,186 65,451,024 196,336,234 OOM OOM 687,089,901 7678 4,806,660 6243
# of best / # of instances 3/15 3/15 15/15
average running time(s) 1347.1 15.7 818.7 617.5
[…]
The comparative results on SET-2 and SET-3 are presented in Tables 3 and 4. The first four columns list the instance information. The remaining columns show the best-known objective value (BKV), the running time of the exact algorithm (time), the mean running time that the heuristic algorithm first hits the BKV (htime), the kernel graph size ($|V_{\text{ker}}|$) and the time of generating a ${}_{\text{ker}}$). In the last two rows, we show the statistical results. Note that the htime of RLS has included the pre-processing time for building graph kernels.
Table 3 shows that RLS outperforms other methods on all SET-2 instances in terms of both best-known objective value and runtime. It is known that RLS proves optimality if all vertices are reduced (i.e. $|V_{kernel}|=0$), hence 11 instances have been solved to optimality by RLS. In contrast, LA-B&B can only solve 1 out of these 47 instances. Moreover, RLS and ALS hit the BKV in just a few seconds.

### v2

@@ 6 Computational Experiments
This section is dedicated to a comprehensive evaluation of the proposed reduction rules and heuristic approaches. To evaluate the algorithm performance, we compare both solution quality and runtime between the RLS and other benchmark algorithms. To evaluate the effect of the reduction rules, we compare the algorithms with and without reductions; we also conduct an ablation study to evaluate the effect of each reduction rule.
@@ 6.1 Datasets
The proposed reduction rules and the RLS algorithm are evaluated on three sets of $216+47+15=278$ instances, covering various application scenarios.
SET-1. The first set of instances consists of $216$ standard instances with up to $400$ vertices and up to $71,820$ edges. These instances were first introduced by Colombi et al. (2017) and tested in Nogueira et al. (2021), Hosseinian and Butenko (2019), Zheng et al. (2024). This set of instances consists of 12 graphs with 125 to 400 vertices and 6,963 to 71,820 edges (available at https://or-dii.unibs.it/instances/Instances_GISP.zip).
@@
SET-3. To further evaluate the algorithms on large instances and diverse real-world scenarios, we introduce 15 instances with up to 18 million vertices and 261 million edges. The first 4 smaller instances represent ecosystem and road/facility networks, while the remaining 11 larger instances are drawn from hyperlinks, citations, and social networks. The instances were generated using the same method for generating SET-2 (available at https://github.com/PlutoAiyi/RLS/tree/main).
@@ 6.2 Experimental Settings
All experiments are conducted on a machine with an Intel(R) Xeon(R) Platinum 8360Y CPU @ 2.40GHz and 2TB RAM.
Benchmark Algorithms. To evaluate the effectiveness of our proposed reduction rules and the RLS algorithm, we compare them with the best-performing algorithms for the GIS problem.
@@
  • CPLEX: The mixed integer programming solver with version 22.12 (released on December 2024).
@@
  • ALS: The adaptive local search algorithm proposed by Zheng et al. (2024)
@@
The source codes of CB&B, LA-B&B, ILS-VND and ALS were made available by the authors. All source codes are available at https://github.com/PlutoAiyi/RLS.
Parameter Settings. The RLS algorithm requires two parameters: the search tolerance $L$ and the perturbation parameter $\epsilon\in(0,1)$. We set $L=10|V|$ where $|V|$ is the number of vertices of the input graph and $\epsilon=0.2$. This setting was obtained using the general IRACE automatic parameter configuration tool (López-Ibáñez et al. 2016). For large instances like soc-pokec, LiveJ and uk2002, we decrease the search tolerance as $L=0.5|V|$ to encourage explorations within the cutoff time.
@@ 6.3 Performance Comparison with Existing Algorithms
We first compared five algorithms on the SET-1 instances. The two branch & bound algorithms CB&B and LA-B&B ran with a cutoff time of 3 hours (10,800 seconds) for each instance. If an exact algorithm fails to finish within the cutoff time, then we report the best solution found so far. The remaining three heuristic algorithms ran 10 trials with a cutoff time of 30 seconds for each instance. We reported the largest value among the best-known objective values obtained in these trials. The setting is consistent with that in (Hosseinian and Butenko 2019, Nogueira et al. 2021).
In Table 2, we summarize the computational results on SET-1. In the last two columns, we report the number of instances where each algorithm achieves the best (#Best) and worst (#Worst) performance, with respect to the best-known objective value and runtime, respectively. Note that the runtime of the RLS refers to the wall-clock time, including both pre-processing and local search. We observe that, in all instances, the best-known objective values obtained by local search algorithms are comparable to those obtained by exact algorithms. In particular, if an exact algorithm obtains the optimal solution within 3 hours, local search algorithms are also able to find an optimal solution. However, when neither of the exact algorithms finds the optimal solution, local search algorithms always produce a better solution. Besides, the exact LA-B&B dominates CB&B whereas the heuristics ALS and the RLS dominate ILS-VND in terms of both solution quality and runtime. Therefore, we only consider LA-B&B, ALS, and the RLS for even larger instances in SET-2 and SET-3. Detailed results on SET-1 are shown in Table LABEL:tab:_Set1_detail in Appendix B.
@@
bio-yeast 1,458 980 968 68,574 1,111 <<1 66,545 1,062 2 68,572 1,113 9 68,574 1,111 <<1 0 <<1
soc-wiki-Vote 889 727 2,187 37,358 648 <<1 35,973 615 <<1 37,358 648 8 37,358 648 <<1 0 <<1
vc-exact_131 2,980 1,191 4,169 109,641 2,098 <<1 107,812 2,045 14 109,392 2,121 14 109,641 2,098 <<1 0 <<1
web-edu 3,031 4,813 1,661 114,140 1,752 <<1 109,856 1,644 10 113,781 1,746 16 114,140 1,753 <<1 98 <<1
MANN_a81 3,321 3,118 3,362 117,327 2,084 <<1 114,616 1,999 16 116,907 2,125 14 117,327 2,084 <<1 0 <<1
tech-routers-rf 2,113 1,539 5,093 97,799 1,668 <<1 93,690 1,588 54 97,799 1,668 11 97,799 1,667 <<1 0 <<1
vc-exact_039 6,795 2,574 8,046 259,542 5,018 <<1 254,941 4,908 253 258,137 5,164 9 259,542 5,018 <<1 0 <<1
ca-GrQc 4,158 3,179 10,243 174,782 2,991 9 167,846 2,871 45 174,654 2,997 15 174,782 2,994 11 282 <<1
vc-exact_038 786 10,321 3,703 12,015 208 589 10,406 168 1,994 12,014 208 12 12,015 207 4 697 <<1
vc-exact_078 11,349 4,289 13,450 438,671 8,330 <<1 431,371 8,128 1,074 435,412 8,671 <<1 438,671 8,330 <<1 0 <<1
vc-exact_087 13,590 5,126 16,114 518,500 10,025 2 509,734 9,809 2,319 517,500 10,083 2 518,500 10,026 <<1 0 <<1
vc-exact_107 6,402 15,646 5,594 407,878 6,763 <<1 396,456 6,488 10,646 400,223 6,846 1 407,878 6,763 <<1 0 <<1
vc-exact_151 15,783 12,144 12,519 530,138 9,526 2 516,445 9,135 2,402 522,462 9,764 1 530,138 9,526 <<1 0 <<1
vc-exact_167 15,783 18,235 6,428 465,178 7,797 1 449,579 7,438 10,469 457,272 7,923 2 465,178 7,797 <<1 0 <<1
bio-dmela 7,393 18,897 6,672 302,992 4,952 <<1 280,897 4,474 170 302,449 4,959 10 302,992 4,951 <<1 44 <<1
vc-exact_011 9,877 19,102 6,871 303,793 5,369 1 286,137 4,950 4,818 301,789 5,487 7 303,793 5,387 <<1 172 <<1
web-spam 4,767 8,089 28,386 190,507 3,285 4.61 180,446 3,130 53 190,344 3,299 14 190,459 3,289 13 487 <<1
vc-exact_026 6,140 27,159 9,608 201,509 3,887 1 177,902 3,372 668 201,316 3,922 17 201,503 3,900 <<1 300 <<1
vc-exact_001 6,160 19,655 20,552 215,941 4,129 2 194,984 3,709 9,575 215,668 4,175 13 215,941 4,128 <<1 134 <<1
vc-exact_024 7,620 11,367 35,926 236,109 4,632 10,783 226,540 4,438 244 235,440 4,735 14 236,018 4,659 12 1,255 <<1
C1000.9 1,000 11,989 37,432 5,263 99 10,800 6,104 116 430 7,229 126 12 7,239 125 10 1,000 <<1
tech-WHOIS 7,476 42,181 14,762 325,584 5,368 5 309,903 5,050 218 325,327 5,367 14 325,581 5,371 13 338 <<1
p-hat700-3 700 30,398 31,242 4,025 68 10,800 3,030 52 7,436 4,473 78 <<1 4,473 78 <<1 700 <<1
vc-exact_008 7,537 35,678 37,155 247,091 4,800 3 210,126 4,089 9,093 246,771 4,839 14 247,091 4,810 <<1 574 <<1
keller5 776 17,800 56,910 2,373 58 10,800 2,860 54 6,327 3,877 69 1 3,877 69 2 776 <<1
vc-exact_194 1,150 39,520 41,331 3,734 77 10,708 4,386 73 1,324 5,419 86 <<1 5,419 86 <<1 1,150 <<1
hamming10-4 1,024 66,369 23,231 3,256 52 8,108 3,073 45 2,069 3,660 54 <<1 3,660 54 <<1 1,024 <<1
brock800_2 800 26,629 84,805 1,177 49 10,800 1,996 36 10 2,473 44 <<1 2,473 43 <<1 800 <<1
brock800_4 800 26,831 85,126 1,583 48 10,800 1,979 39 7,215 2,501 43 <<1 2,501 42 <<1 800 <<1
ca-HepPh 11,204 57,721 59,898 372,096 6,094 10,800 348,738 5571 422 370,779 6,083 6 372,079 6,098 17 2,165 <<1
p-hat700-2 700 90,788 32,134 2,870 41 10,140 2,345 37 2,078 2,971 42 <<1 2,971 42 <<1 700 <<1
vc-exact_196 1,534 93,273 32,809 3,916 62 10,800 4,273 57 66 5,321 76 <<1 5,321 76 <<1 1,534 <<1
p-hat700-1 700 135,804 47,847 730 11 10,800 844 11 5 844 11 <<1 844 11 <<1 700 <<1
ca-AstroPh 17,903 47,135 149,837 604,288 10,468 10,800 558,736 9,752 2,544 602,320 10,521 2 604,019 10,491 14 4,985 1
C2000.9 2,000 147,548 51,920 3,958 65 9,489 5,477 73 11 6,396 86 3 6,396 86 3 2,000 <<1
DSJC1000.5 1,000 59,694 189,980 925 39 7,645 1,515 30 694 1,861 32 1 1,861 32 <<1 1,000 <<1
socfb-MIT 6,402 123,064 128,166 7,098 124 10,800 110,677 1,779 20 129,807 2,142 17 129,917 2,132 16 5,494 <<1
p-hat1500-3 1,500 66,206 210,800 4,097 132 10,800 3,877 83 704 7,501 148 14 7,501 148 16 1,500 <<1
socfb-UCSB37 14,917 356,918 125,297 16,708 287 10,800 219,965 3,286 256 254,965 4,013 16 255,924 3,994 22 13,655 3
socfb-Duke14 9,885 374,731 131,706 11,565 198 10,800 125,940 1,887 39 151,169 2,415 15 151,365 2,390 18 9,003 4
p-hat1500-2 1,500 411,372 143,918 3,542 53 10,800 2,611 43 2,150 4,253 68 <<1 4,253 68 <<1 1,500 <<1
socfb-Stanford3 11,586 136,823 431,486 7,250 122 10,800 215,381 3,932 265 251,350 4,441 12 251,932 4,436 14 9,380 2
socfb-UConn 17,206 448,038 156,829 16,996 279 10,800 231,718 3,471 322 274,637 4,329 12 275,320 4,299 18 16,025 5
p-hat1500-1 1,500 411,744 427,583 0 0 10,800 1,184 18 5,499 1,287 19 <<1 1,287 19 1 1,500 <<1
C2000.5 2,000 489,894 509,270 0 0 10,800 1,581 24 865 1,849 26 2 1,849 26 6 2,000 1
keller6 3,361 759,816 266,766 0 0 10,800 3,545 55 1,561 5,236 73 10 5,236 73 16 3,361 <<1
C4000.5 4,000 1,957,884 2,039,848 0 0 10,800 1,611 25 1,659 1,948 28 20 1,952 28 22 4,000 9
# of best / # of instances 24/47 1/47 18/47 41/47
average running time(s) 5,402.1 2,087.4 7.9 5.9 1.4
Next, we compared the RLS with LA-B&B, ALS and CPLEX on SET-2 and SET-3. Each of the two heuristic ALS and the RLS were run 10 trials and reported its best known objective value. Each trial was given a cutoff time of 30 seconds. In total, each heuristic algorithm ran 300 seconds per instance in SET-2. The settings are the same as in Zheng et al. (2024). The cutoff time of exact algorithms was directly set to 3 hours (10,800 seconds) per instance in SET-2. However, for SET-3, the heuristic trial time was increased to 200 seconds. For large instances like soc-pokec and LiveJ, the cutoff time of ALS and the RLS were given 1,000 seconds per trial. For the larger uk2002, the cutoff time of LA-B&B and CPLEX was 100,000 seconds, and the cutoff time of ALS and the RLS is 10,000 seconds per trial. In particular, for the three instances soc-pokec, LiveJ and uk2002, the RLS excluded $R6$ and R12R121414, as these reductions exhibited considerably slower performance compared to the other rules in the experiments. For LA-B&B and CPLEX, if the algorithm did not finish within the total running time, the best solution found so far was reported.
[…]
@@
foodweb-wet 128 1,509 566 2,464 44 <<1 2,464 44 8 2,464 44 <<1 2,464 44 <<1 125 <<1
foodweb-dry 128 1,571 535 2,226 42 <<1 2,226 42 8 2,226 42 <<1 2,226 42 <<1 125 <<1
USAir97 332 1,589 537 10,308 194 <<1 9,507 178 407 10,308 194 <<1 10,308 194 <<1 0 <<1
powergrid 4,941 1,727 4,867 193,339 3,710 <<1 190,537 3,684 108 193,095 3,754 59 193,339 3,537 <<1 0 <<1
CondMat 23,133 46,769 46,670 711,517 12,821 12 685,442 12,061 5,749 705,901 1,306 14 711,334 12,870 45 1,648 <<1
Email 265,009 91,093 273,388 13,054,427 257,435 118 OOM OOM 13,054,427 257,394 1 0 1
Epinion 75,879 202,653 203,087 3,052,059 57,976 60 OOM 3,038,533 58,267 18 3,052,059 58,008 1 253 1
Dblp 317,080 787,268 262,598 9,737,911 171,421 790 OOM OOM 9,734,653 171,403 64 11,851 11
cnr-2000 325,557 685,686 2,053,283 13,367,460 260,331 10,800 OOM OOM 13,368,670 260,716 482 59,052 399
WikiTalk 2,394,385 3,494,873 1,164,692 118,708,326 2,342,835 3,566 OOM OOM 118,708,326 2,342,861 333 8 333
BerkStan 685,230 1,662,833 4,986,637 26,036,210 501,197 10,800 OOM OOM 26,031,704 501,410 1,112 114,191 1,031
As-Skitter 1,696,415 5,550,661 5,544,637 66,603,360 1,264,844 10,800 OOM OOM 66,582,254 1,265,266 723 31,173 636
soc-pokec 1,632,803 11,152,945 11,149,019 4,146,172 84,217 10,800 OOM OOM 43,999,402 833,989 934 881,319 134
LiveJ 4,846,609 10,710,579 32,140,658 0 0 10,800 OOM OOM 168,431,056 3,258,831 904 401,307 469
uk2002 18,483,186 65,451,024 196,336,234 0 0 100,000 OOM OOM 687,089,901 13,716,043 7,678 4,806,660 6,243
# of best / # of instances 11/15 3/15 3/15 11/15
average running time(s) 10,570.1 1,256 15.7 818.7 617.5
Table 3 shows that the RLS outperforms other methods on all SET-2 instances in terms of both best-known objective value and runtime. It is known that the RLS proves optimality if all vertices are reduced (i.e., $|V_{kernel}|=0$), hence 11 out of 47 instances have been solved to optimality by the RLS. In contrast, LA-B&B can only solve 1 out of these 47 instances. Similarly, CPLEX fails to solve instances with 40,000 variables (e.g., vc-exact_024 and C1000.9). Moreover, CPLEX even fails to report a feasible solution for 4 large instances. In contrast, the RLS and ALS hit the BKV with an average hitting time of no more than 8 seconds.
Table 4 compares algorithms on SET-3 instances. OOM indicates that the algorithm runs out of memory. Both ALS and LA-B&B quickly become infeasible as instance size increases. CPLEX again fails to report a feasible solution for large instances. In contrast, the RLS computes feasible solutions even for large instances such as uk2002 . For instances where the benchmarks are feasible, the RLS consistently achieves the best net benefit. For the infeasible instance Email, the RLS even finds the optimal solution.
6.4 Evaluation of Reduction Rules
To evaluate the effectiveness of the reduction rules, we conducted the following two experiments.